OpenStudy (geekfromthefutur):

Ismael is comparing cell phone plans before upgrading his phone. Ameri-Mobile offers a low activation fee, but a high monthly payment. Cell-U-Later offers a lower monthly rate, but the activation fee is higher. Create a possible algebraic expression for both Ameri-Mobile and Cell-U-Later that shows the amount paid after an unknown amount of months have passed. Justify how you created those expressions and identify what each term and factor represents in terms of the cell phone plans.

OpenStudy (geekfromthefutur):

I have been stuck on this math question for 30 minutes i really could use some help with this. Plz i will become fan and give medal.

OpenStudy (geekfromthefutur):

@jdoe0001

OpenStudy (anonymous):

make up some numbers say America Mobile charges $100 activation and $30 a month Cell U Later $140 activation but only $20 a month you can make up your own

OpenStudy (anonymous):

then if say x is the number of months after you get the plan, the cost of AM will be 100+30x whereas the cost of CUL is 140+20x

OpenStudy (anonymous):

good?

OpenStudy (jdoe0001):

geekfromthefutur do you have cable tv?

OpenStudy (geekfromthefutur):

Im understanding it

OpenStudy (geekfromthefutur):

Direct tv

OpenStudy (anonymous):

ok pm me if u need a better understanding.

OpenStudy (jdoe0001):

ok well... check the original bill of DirectTV for the DirectTV guy to just show up and SETUP the antenna and service and put the settop box, there's a charge now the DirectTV guy isn't at your home every month he did that only ONCE, a one-time charge after that you have to pay some fixed cost for the services you requested is pretty much the same amount month by month so say the DirectTV charges applied to your case will be f(x) = (INITIAL SETUP FEE THE GUY DID) + (MONTHLY RATE) for 1,2, 3, 4 months will be f(x) = (INITIAL SETUP FEE THE GUY DID) + (MONTHLY RATE) *1 f(x) = (INITIAL SETUP FEE THE GUY DID) + (MONTHLY RATE) *2 f(x) = (INITIAL SETUP FEE THE GUY DID) + (MONTHLY RATE) *3 f(x) = (INITIAL SETUP FEE THE GUY DID) + (MONTHLY RATE) *4 for "x" months will be f(x) = (INITIAL SETUP FEE THE GUY DID) + (MONTHLY RATE) *x

OpenStudy (jdoe0001):

many service contracts do the same scheme to start up the service, they need to setup a new client, add the services, synchronize the billing to that new client so that's expense the company takes, and thus charges the client, but only ONCE they don't do that every month, just ONE TIME and then just go along with the monthly charges usually will be like f(x) = (SOME VALUE HERE) + (OTHER VALUE) x

OpenStudy (geekfromthefutur):

ok yea now i understand lol thx dude ur a life saver

OpenStudy (jdoe0001):

yw